OpenStudy (anonymous):

Use a proportion to solve this problem. Jim drove 250 miles in 5 hours. If he can keep the same pace, how long will it take him to drive 1,100 miles ? HELP please !

OpenStudy (scorcher219396):

If you know that two different qualities belong to the same event, and you're trying to find a missing quality from another event (you are) Then you'd set it up like this \[\frac{ 250 miles }{ 5 hours }=\frac{ 1100 miles }{ x hours }\] Now cross multiply and divide to solve for x 250x=5500 /250 x=22 hours

OpenStudy (scorcher219396):

A more lengthy explanation, (ignore this completely if it confuses you,) A formula for distance, rate and time is d=rt If you divide both sides by t, then the new equation is d/t=r Then you can plug in the information you're given, 250miles/5hours, which if you were to simplify that would give you your rate, but you don't need to do that Since the rate is the same as the if he were going 1100 miles, you can set that (250miles/5hours) equal to 1100 miles/x hours because you're basically just using the d=rt formula twice, and substituting